OpenStudy (xishem):

At elevated temperatures, nitrous oxide decomposes according to the equation given below...

OpenStudy (xishem):

\[2N_2O(g) \rightarrow 2N_2(g)+O_2(g)\]Given the following data, plot the appropriate graphs, and determine whether the reaction is first order or second order. What is the value of the rate constant for the consumption of N_2O?

OpenStudy (xishem):

I was able to find that the reaction is first order, and the equation for that graph is\[y=-0.0023x-1.3863\]Which would suggest that the rate constant for this reaction is\[0.0023(\min)^{-1}*\frac{60s^{-1}}{1\min^{-1}}=0.138s^{-1}\]However, this doesn't match the answer in the back of the back which is:\[3.79*10^{-5}s^{-1}\]Any help?

OpenStudy (turingtest):

you just got your conversion upside-down by trying to let things be inverse\[0.0023(\min)^{-1}*\frac{1\min}{60s}=3.83\times10^{-5}s^{-1}\]ok it's not exact, so you must be a little off somewhere else. A lot closer though :D

OpenStudy (xishem):

Ah! Thank you. It ended up being an issue with excel rounding the coefficient behind t^1 to -0.0023, whereas the value with many more sigfigs is -0.0022764840. Using this new value instead of the rounded value produced the correct result of 3.79 * 10^-5. And what is the reason behind inverse factors not working? Is it because I was assuming this...\[\frac{60*(s)^{-1}}{1 * (\min)^{-1}}\]When I should have been assuming this?\[\frac{(60*s)^{-1}}{(1*\min)^{-1}}\]

OpenStudy (turingtest):

yes, that would work would the 60^-1, because that way you divide by 60 (which is what you need to have happen), though I don't do it that way personally.